Three years back, a father was 24 years older than his son. At present the father is 5 times as old as the son. How old will the son be three years from now?

Found 2 solutions by math_tutor2020, ikleyn:
Answer by math_tutor2020(3817)


x = Son's age right now
5x = Father's age right now
3 years agoPresent3 years in the future
Son’s Agex-3xx+3
Father’s Age5x-35x5x+3


3 years ago, the father was 24 years older than his son
father's past age = (son's past age) + 24
5x-3 = (x-3) + 24

I'll let the student solve for x.
Once you determine the value of x, compute x+3 to get the final answer.

Answer by ikleyn(52788)

.
Three years back, a father was 24 years older than his son. At present the father is 5 times as old as the son.
How old will the son be three years from now?
~~~~~~~~~~~~~~~~~~~~~~~

Since Three years back the father was 24 years older than his son,

at present we have the same difference: today the father is 24 years older than his son.


So, two statements are valid simultaneously:

    (1)  today the father is 24 years older than his son,

and

    (2)  At present, the father is 5 times as old as the son.


Let the son age is x years; then the father is 5x years old, and

    5x - x = 24.


So, 4x = 24,  x = 24/4 = 6 years,  so the son is 6 years old now and the father is 6+24 = 30 years old now.


In 3 years from now, the son will be  6 + 3 = 9 years old.    ANSWER

Solved.